OpenStudy (anonymous):

Part A: Max rented a motorbike at $465 for 5 days. If he rents the same motorbike for a week, he has to pay a total rent of $625. Write an equation in the standard form to represent the total rent (y) that Max has to pay for renting the motorbike for x days. (4 points) Part B: Write the equation obtained in Part A using function notation. (2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (anonymous):

i just need help with part c

OpenStudy (anonymous):

@KryoWolf what did you get for part A and B?

OpenStudy (anonymous):

Part A: 625-465=$160 he has to pay for 2 days=$160 rent paid for 1 day=260/2=$80 amount paid for 5days=8*5=$400 But he paid $465 Difference 465-400=$65 so every one has to pay $65+rent for the no. of days total rent for x days y=65+80 x (dollars) Part B: f(x)=65+80 x

OpenStudy (anonymous):

this

OpenStudy (anonymous):

i dont know how to graph it though

OpenStudy (anonymous):

The x- axis will show the number of days and the y-axis will show the total cost. I know how to graph it but I don't know how to derive the formula. Good team work mate! choose x-values such as 1, 2, 3, 4, 5, 6 and plug then in the formula and each will represent a point on the graph. Notice here that 65 here is the y-intercept.

OpenStudy (anonymous):

thats the problem i dont know how to do the formula either

OpenStudy (anonymous):

@Boblovesmath

OpenStudy (anonymous):

know anyone who does?

OpenStudy (anonymous):

I think the formula is correct, just use x-values to represent the days and the output will represent the cost. So for example if you plug in 1 in the formula, you would get 80 + 65 = 145 $ which is the cost of one night. then 80 + 130 = 210, which is the cost of the second night and so on. After plugging in all the points you should notice that this is a linear function.

OpenStudy (anonymous):

ohh ok thanks

OpenStudy (anonymous):

@Boblovesmath its 1.69

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

my pleasure :)

OpenStudy (anonymous):

i is now ur fan

OpenStudy (anonymous):

@KryoWolf me 2 :)